# Bistable quartic soliton in saturable nonlinear media

### v1

@@ Abstract
In this letter, for the first time, to the best of our knowledge, we theoretically demonstrate the existence of novel bistable quartic soliton (BQS) in saturable nonlinear media. We propose a realistic dispersion-engineered ridge waveguide based on Lithium Niobate $(\text{LiNbO}_{3})$ that offers a suitable environment to excite the family of BQS. Adopting the variational method we analytically establish the amplitude-width relation which reveals that stable QSs with identical duration but two different amplitudes can coexist. We further investigate the robustness of such BQS under perturbation by performing the linear stability analysis.
preprint: AIP/123-QED
@@
An optical soliton is a stable wave that retains its shape during long-distance propagation, achieved through the balance of group-velocity dispersion (GVD) and Kerr nonlinearity [1]. In addition to traditional solitons, a new class known as quartic soliton (QS) was introduced in early 90s [2, 3, 4, 5]. QS emerges in a specific dispersion regime where the coefficients of 3rd-order dispersion ($\beta_{3}$) is zero, and both 2nd ($\beta_{2}$) and $4^{th}$ ($\beta_{2}$) order dispersion are negative [6]. This unique point on the dispersion curve is referred to as the quartic point and essentially the inspiration to the author to coin the term quartic soliton for the first time [7]. This new soliton type differs from conventional Kerr-solitons by having a secant hyperbolic square [3] pulse shape and may exhibit radiationless oscillatory tails contingent on its propagation constant [5].
@@
After its first theoretical introduction in the 90s, research on QS has largely stagnated due to the adverse effects of Raman-induced frequency redshift. This redshift can cause the soliton frequency to shift away from the quartic point, where QS becomes non-existent. Notably, the Raman-induced frequency redshift ($\Delta\omega_{R}$) is significantly impacted by pulse duration ($t_{0}$), scaling as $\Delta\omega_{R}\propto t^{-6}_{0}$ for QS, compared to the conventional soliton’s scaling of $\Delta\omega_{R}\propto t^{-4}_{0}$ for short pulses [8] in optical fibers. Achieving the desired dispersion profile for QS has been a significant challenge; however, advancements in nanotechnology for waveguide fabrication have enabled the customization of dispersion characteristics. Exploiting such technique, recent theoretical and experimental studies have focused [9, 10, 11, 12, 13] specifically on the sub-branch of QS which is formally known as pure quartic soliton (PQS)[9, 14]. PQSs are generated through the interaction of negative quartic dispersion and Kerr nonlinearity under conditions of zero GVD. Unlike Kerr solitons, which follow the energy-width relationship $E\propto t^{-1}_{0}$, PQSs exhibit a relationship of $E\propto t^{-3}_{0}$ [11], indicating they can transport more energy for the same pulse duration. This characteristic positions PQSs as promising candidates for high-power laser applications[13].
@@
Though the investigation of QS has been studied in various modulated dispersion environments [15, 16] its exploration is limited within the domain of Kerr nonlinearity. Perhaps the reason behind this is, QSs, PQSs, and conventional Kerr solitons are categorized as members of the same family generalized-dispersion Kerr solitons [17]. This limitation raises interest in exploring the existence of QS in broader types of nonlinear systems. Motivated by this idea, in this letter, we focus on investigating the formation of QS in an optical medium offering saturable nonlinearity (SN) which, to the best of our knowledge, has never been explored before.
@@
The SN is typically associated with materials such as photo-refractive (PR)[18, 19, 20] crystals and semiconductor doped glass (SDG) [21]. In this analysis, we focus on Lithium Niobate (LiNbO3, LN), recognized for its significant PR properties, including a high $\chi^{(3)}$ nonlinearity of $n_{2}=2.5\times 10^{-19}$ m2/W and a broad optical transparency range (350 nm - 4500 nm)[22]. The PR nonlinearity in LiNbO3 arises from the light-induced space charge field $E_{sc}$, which modulates the refractive index through the electro-optic effect [23], contingent upon appropriate bias and orientation of the sample. Experimental evidence demonstrates that self-focusing SN can occur in LiNbO3 under an external electric field aligned with the optic axis of the crystal[19]. Notably, this PR effect is carrier-dependent, resulting in a delayed response compared to the Kerr effect when altering the refractive index. Effective modulation in PR materials is achieved through the application of high-irradiance femtosecond laser pulses, which facilitate the generation of SN[24, 25, 26].
@@
In this letter, we investigate the existence of bistable quartic-solitons (BQSs) within a LiNbO3 based waveguide that exhibits SN. Our theoretical analysis demonstrates that these robust BQSs can arise in suitable dispersion conditions. We examine the temporal dynamics of the BQSs under shock effects, which act as perturbations specially for SDGs. Additionally, we assess the stability of these BQSs against input amplitude noise and establish a stability map across a parametric space. Our findings indicate a pathway to realizing a new class of solitary waves in mediums with SN, which could be significant for applications in fiber lasers and optical communications.
@@
We design a realistic ridge waveguide composed of LiNbO3, featuring external bias along a transverse direction. The GVD profile for a quasi-transverse-electric (TE) mode of the extraordinary ray is illustrated in Fig. 1. To excite SN, an external electric-field bias is applied along the optic axis[19], with uniform irradiation to avoid oversaturation[18]. The waveguide cross-section, system set-up, and mode confinement are illustrated as an inset (upper and lower) in Fig. 1. The GVD-profile exhibits the quartic point (red dot) encircled by a shaded elliptical region where the approximation $\beta_{3}\approx 0$ is valid and the GVD profile can be expressed by including only $\beta_{2}$ and $\beta_{4}$, terms respectively. In the proximity of the quartic point, GVD is anomalous ($\beta_{2}<0$) and GVD curvature is negative ($\beta_{4}<0$). Exactly at the quartic point, for operating wavelength $\lambda_{0}\approx 1.73\mu$m the dispersion coefficients are calculated as $\beta_{2}=-0.014$ ps2/m, $\beta_{3}=2.51\times 10^{-5}$ ps3/m and $\beta_{4}=-1.30\times 10^{-5}$ ps4/m, respectively. Note that, for focusing nonlinearity, these are the primary conditions for the exploration of localized QS and we carefully design the waveguide structure to achieve the correct sign and curvature of the GVD at operating wavelength. The optical pulse propagation in the waveguides where the nonlinear response saturates beyond a threshold power can be modeled by standard NLSE. Under the slowly varying envelope approximation, the complex-valued electric field satisfies the following nonlinear Schrödinger equation (NLSE) [1],
$i\partial_{\xi}\psi+\sum_{m\geq 2}i^{m}\delta_{m}(\partial_{\tau})^{m}\psi+(1+i\tau_{sh}\partial_{\tau})f(|\psi|^{2})\psi=0,$ (1)
@@
here, we perform the typical scaling $\psi\rightarrow A/\sqrt{P_{0}}$, $\tau\rightarrow(t-v_{g}/z)/t_{0}$, $\xi\rightarrow z/L_{D}$, where $L_{D}=t^{2}_{0}/|\beta_{2}|$ is the dispersion length, $\delta_{m}=\beta_{m}/(m!|\beta_{2}|t^{m-2}_{0})$, $\beta_{m}$ is the mth order dispersion coefficient at the carrier frequency $\omega_{0}$. The normalized self-steepening parameter, relevant for SDG fibers, is defined as, $\tau_{sh}=(\omega_{0}t_{0})^{-1}$ where $t_{0}$, $P_{0}$, $v_{g}$, $\gamma$ being the input pulse duration, input peak power, group velocity and nonlinear coefficient, respectively. For our proposed waveguide, the refractive index of the LN core at the operating wavelength ($\lambda_{0}\approx 1.73$ $\mu$m) is around $n_{core}=2.13$ which offers a large refractive index contrast between core and silica cladding ($n_{clad}=1.44$), thereby resulting a tight mode confinement as indicated in the inset Fig. 1. For the input pulse width $t_{0}=55$ fs, the normalized dispersion coefficients are calculated as $\delta_{4}=-0.0125$, and $\delta_{3}=0.0053$. Note, the value of $\delta_{3}$ is one order less than $\delta_{4}$. For our system, we adopt the frequently used mathematical form of the SN response [27, 28, 29, 30] as,
@@
Here the coefficient $\delta_{2}$ and $\delta_{4}$ are dominating and we neglect all higher order dispersion terms ($\delta_{m>4}=0$) including the 3OD (as $|\delta_{3}/\delta_{4}|<<1$). For PR crystals, we ignore the shock effect ($\tau_{sh}=0$) and the impact of two-photon absorption is minimal, as the energy of two-photon quanta ($\approx 0.7$ eV) is significantly lower than the energy band gap of LN ($\approx 3.9$ eV) [31]. Additionally, since experiments show no notable Raman shift [32, 33] in LN, this effect is also excluded.The SN media can excite solitary waves in the form $\psi(\xi,\tau)=\sqrt{\psi_{s}(\tau)}e^{iq\xi}$ , where $q$ stands for propagation constant [27]. In our case, however, we seek for a stationary solution in the form $\psi(\xi,\tau)=g(\tau)e^{iq\xi}$ and substituting it in Eq. (3) we can have the equation for $g(\tau)$ as,
@@
$\Delta\tau_{w}=\mathcal{A}^{2}\tau_{sh}\xi\digamma_{s},$ (8)
where a hyperbolic secant function is considered as an ansatz and the parameter ϝ$\digamma_{s}$ is defined as, ϝ$\digamma_{s}=\frac{1}{m}-\frac{2}{m}\frac{\ln(\sqrt{m}+\sqrt{m+1})}{m\sqrt{m}\sqrt{m+1}}-\left[\frac{\cos^{-1}(1+2m)}{2m}\right]^{2}$ with $m=s\mathcal{A}^{2}$. In Fig. 2 ($e$)-($f$), the white dotted lines indicate the analytical prediction (based on Eq. (8)) of temporal shift which shows an excellent agreement with the full numerical result.
Stability Analysis:
In this section we investigate the stability of BQS in saturable media through linear stability analysis (LSA). In this approach, the stationary QS, denoted as $g(\tau)$, is subjected to small amplitude perturbations represented by $v(\tau)$ and $w(\tau)$. The optical field is expressed as $\psi(\tau,\xi)=\left[g(\tau)+w^{*}(\tau)e^{h^{*}\xi}+v(\tau)e^{h\xi}\right]e^{iq\xi}$. By substituting this form into the governing equation Eq. (3) and performing linearization with respect to the perturbations, we derive two linear systems for $v$ and $w$ that yield the growth rate $h$. This process culminates in an eigenvalue problem as follows,
@@
The stability phase plot (see Fig. 3) in the $\mathcal{A}$-$s$ parametric space reveals that BQSs are stable within a limited region where the maximum instability eigenvalue, $\Re[h_{max}]$, is zero. To confirm the LSA findings, QS stability was examined by introducing a $10\%$ noise in pulse amplitude at three distinct settings. Utilizing the Crank-Nicolson algorithm, the evolution of the pulse was analyzed for $s=0.5$ and $\delta_{4}=-0.0125$. The results indicated that the QSs were linearly unstable at points 1 ($\mathcal{A}=6$, $s=0.5$) and 3 ($\mathcal{A}=1$, $s=0.5$), whereas they exhibited stability at point 2 ($\mathcal{A}=1.78$, $s=0.5$). The linear-stability spectra revealed that points 1 and 3 contained positive real part eigenvalues, confirming linear instability, in contrast to point 2, where all eigenvalues were imaginary, indicating linear stability. Additionally, in all cases, the continuum eigenvalue edges at $\Im(h)=\pm q$ and the pair of discrete eigenvalues on the imaginary axis indicates the internal modes contributing to shape oscillations in the soliton [36].
In summary, we demonstrate the existence of novel BQS within an optical medium characterized by SN. Through a theoretical analysis employing variational optimization, we find that unique bistable solitonic states with varying amplitude profiles yet identical widths can persist. A realistic waveguide was engineered using the LN crystal, promoting an advantageous dispersion environment for BQS excitation. Linear stability analysis was conducted to confirm the stability of this unique structure against perturbations. Our findings enhance the understanding of self-organized temporal structures as they develop in saturable nonlinear media, indicating potential applications in high-power laser systems and communication technologies.
T.D. and S.R. was financially supported by IIT Kharagpur, the Ministry of Education of the Government of India, and A.P.L. are supported by Anusandhan National Research Foundation(ANRF) erstwhile SERB, Government of India.
Author Declarations
Conflict of Interest
The authors have no conflicts to disclose.
Author Contributions
Tiyas Das: Conceptualization (equal); Data curation (equal); Formal analysis (equal); Investigation (equal); Methodology (equal); Software (equal); Visualization (equal); Writing – original draft (equal); Writing – review and editing (equal). Samudra Roy: Conceptualization (equal); Data curation (equal); Formal analysis (equal); Investigation (equal); Methodology (equal); Software (supporting); Visualization (equal); Writing – original draft (equal); Writing – review and editing (equal). Anuj Pratim Lara: Conceptualization (supporting); Data curation (equal); Formal analysis (equal); Investigation (equal); Methodology (equal); Software (equal); Visualization (equal); Writing – original draft (supporting); Writing – review and editing (equal)
Data Availability
The data that support the findings of this study are available within the article .

### v2

@@ Abstract
This report presents a theoretical demonstration of a novel bistable quartic soliton (BQS) in saturable nonlinear media, specifically within a realistic dispersion-engineered ridge waveguide made of Lithium Niobate ($\text{LiNbO}_{3}$). The study employs the variational method to establish the amplitude-width relationship, indicating the coexistence of stable solitons with the same duration but differing amplitudes. Additionally, the impact of shock on the bistable soliton is examined through perturbative variational analysis, supported by numerical results. The robustness of the BQS under perturbations is further investigated via linear stability analysis.
@@
An optical soliton is a stable wave that retains its shape during long-distance propagation, achieved through the balance of group-velocity dispersion (GVD) and Kerr nonlinearity [1]. In addition to traditional solitons, a new class known as quartic soliton (QS) was introduced in early 90s [2, 3, 4, 5]. QS emerges in a specific dispersion regime where the coefficients of 3rd-order dispersion ($\beta_{3}$) is zero, and both 2nd ($\beta_{2}$) and $4^{th}$ ($\beta_{2}$) order dispersion are negative [6]. This unique point on the dispersion curve is referred to as the quartic point and essentially the inspiration to the author to coin the term quartic soliton [7]. This new soliton type differs from conventional Kerr-solitons by having a secant hyperbolic square [3] pulse shape and may exhibit radiationless oscillatory tails contingent on its propagation constant [5].
@@
After its first theoretical introduction in the 90s, research on QS has largely stagnated due to the adverse effects of Raman-induced frequency redshift. This redshift can cause the soliton frequency to shift away from the quartic point, where QS becomes non-existent. Notably, the Raman-induced frequency redshift ($\Delta\omega_{R}$) is significantly impacted by pulse duration ($t_{0}$), scaling as $\Delta\omega_{R}\propto t^{-6}_{0}$ for QS, compared to the conventional soliton’s scaling of $\Delta\omega_{R}\propto t^{-4}_{0}$ for short pulses in optical fibers [8]. Achieving the desired dispersion profile for QS has been a significant challenge; however, advancements in nanotechnology for waveguide fabrication have enabled the customization of dispersion characteristics. Exploiting such technique, recent theoretical and experimental studies have focused [9, 10, 11, 12, 13, 14] specifically on the sub-branch of QS which is formally known as pure quartic soliton (PQS) [9, 15]. PQSs are generated through the interaction of negative quartic dispersion and Kerr nonlinearity under conditions of zero GVD. Unlike Kerr solitons, which follow the energy-width relationship $E\propto t^{-1}_{0}$, PQSs exhibit an advantageous energy scaling of $E\propto t^{-3}_{0}$ [11], indicating they can transport more energy for the low pulse duration [14]. This characteristic positions PQSs as promising candidates for high-power laser applications [13].
@@
Though the investigation of QS has been studied in various modulated dispersion environments [16, 17, 18] its exploration is limited within the domain of Kerr nonlinearity. Perhaps the reason behind this is, QSs, PQSs, and conventional Kerr solitons are categorized as members of the same family generalized-dispersion Kerr solitons [19]. This limitation raises interest in exploring the existence of QS in broader types of nonlinear systems. Motivated by this idea, in this report, we focus on investigating the formation of QS in an optical medium offering saturable nonlinearity (SN) which, to the best of our knowledge, has never been explored before.
@@
The SN is typically associated with materials such as photo-refractive (PR) [20, 21, 22] crystals and semiconductor doped glass (SDG) [23]. In this analysis, we focus on Lithium Niobate (LiNbO3, LN), recognized for its significant PR properties, including a high $\chi^{(3)}$ nonlinearity of $n_{2}=2.5\times 10^{-19}$ m2/W and a broad optical transparency range (350 nm - 4500 nm) [24]. The PR nonlinearity in LiNbO3 arises from the light-induced space charge field $E_{\text{sc}}$, which modulates the refractive index through the electro-optic effect [25], contingent upon appropriate bias and orientation of the sample. Experimental evidence demonstrates that self-focusing SN can occur in LiNbO3 under an external electric field aligned with the optic axis of the crystal [21]. Notably, this PR effect is carrier-dependent, resulting in a delayed response compared to the Kerr effect when altering the refractive index. Effective modulation in PR materials is achieved through the application of high-irradiance femtosecond laser pulses, which facilitate the generation of SN [26, 27, 28].
@@
In this brief report, we investigate the existence of bistable quartic-solitons (BQSs) within a LiNbO3 based waveguide that exhibits SN. Our theoretical analysis demonstrates that these robust BQSs can arise in suitable dispersion conditions. We examine the temporal dynamics of the BQSs under shock effects, which act as perturbations specially for SDGs. Additionally, we assess the stability of these BQSs against input amplitude noise and establish a stability map across a parametric space. Our findings indicate a pathway to realizing a new class of solitary waves in mediums with SN, which could be significant for applications in fiber lasers and optical communications.
@@
We design a realistic ridge waveguide composed of LiNbO3, featuring external bias along a transverse direction. The GVD profile for a quasi-transverse-electric (TE) mode of the extraordinary ray is illustrated in Fig. 1. To excite SN, an external electric-field bias is applied along the optic axis [21], with uniform irradiation to avoid oversaturation [20]. The waveguide cross-section, system set-up, and mode confinement are illustrated as an inset (upper and lower) in Fig. 1. The GVD-profile exhibits the quartic point (red dot) encircled by a shaded elliptical region where the approximation $\beta_{3}\approx 0$ is valid and the GVD profile can be expressed by including only $\beta_{2}$ and $\beta_{4}$, terms respectively. In the proximity of the quartic point, GVD is anomalous ($\beta_{2}<0$) and GVD curvature is negative ($\beta_{4}<0$). Exactly at the quartic point, for operating wavelength $\lambda_{0}\approx 1.73\mu$m the dispersion coefficients are calculated as $\beta_{2}=-0.014$ ps2/m, $\beta_{3}=2.51\times 10^{-5}$ ps3/m and $\beta_{4}=-1.30\times 10^{-5}$ ps4/m, respectively. Note that, for focusing nonlinearity, these are the primary conditions for the exploration of localized QS and we carefully design the waveguide structure to achieve the correct sign and curvature of the GVD at operating wavelength. The optical pulse propagation in the waveguides where the nonlinear response saturates beyond a threshold power can be modeled by the standard NLSE. Under the slowly varying envelope approximation, the complex-valued electric field satisfies the following nonlinear Schrödinger equation (NLSE) [1],
$i\partial_{\xi}\psi+\sum_{m\geq 2}i^{m}\delta_{m}(\partial_{\tau})^{m}\psi+(1+i\tau_{sh}\partial_{\tau})f(|\psi|^{2})\psi=0.$ (1)
@@
Here, we perform the typical scaling $\psi\rightarrow A/\sqrt{P_{0}}$, $\tau\rightarrow(t-v_{g}/z)/t_{0}$, $\xi\rightarrow z/L_{D}$, where $L_{D}=t^{2}_{0}/|\beta_{2}|$ is the dispersion length, $\delta_{m}=\beta_{m}/(m!|\beta_{2}|t^{m-2}_{0})$, $\beta_{m}$ is the mth order dispersion coefficient at the carrier frequency $\omega_{0}$. The normalized self-steepening parameter, relevant for SDG fibers, is defined as, $\tau_{sh}=(\omega_{0}t_{0})^{-1}$ where $t_{0}$, $P_{0}$, $v_{g}$, $\gamma$ being the input pulse duration, input peak power, group velocity and nonlinear coefficient, respectively. For our proposed waveguide, the refractive index of the LN core at the operating wavelength ($\lambda_{0}\approx 1.73$ $\mu$m) is around $n_{\text{core}}=2.13$ which offers a large refractive index contrast between core and silica cladding ($n_{\text{clad}}=1.44$), thereby resulting in tight mode confinement as indicated in the inset Fig. 1. For the input pulse width $t_{0}=55$ fs, the normalized dispersion coefficients are calculated as $\delta_{4}=-0.0125$, and $\delta_{3}=0.0053$. Note, the value of $\delta_{3}$ is one order less than $\delta_{4}$. For our system, we adopt the frequently used mathematical form of the SN response [29, 30, 31, 32] as,
@@
Here the coefficients $\delta_{2}$ and $\delta_{4}$ are dominating and we neglect all higher order dispersion terms ($\delta_{m>4}=0$) including the 3OD (as $|\delta_{3}/\delta_{4}|<<1$). For PR crystals, we ignore the shock effect ($\tau_{sh}=0$) and the impact of two-photon absorption is minimal, as the energy of two-photon quanta ($\approx 0.7$ eV) is significantly lower than the energy band gap of LN ($\approx 3.9$ eV) [33]. Additionally, since experiments show no notable Raman shift [34, 35] in LN, this effect is also excluded.The SN media can excite solitary waves in the form $\psi(\xi,\tau)=\sqrt{\psi_{s}(\tau)}e^{iq\xi}$ , where $q$ stands for propagation constant [29]. In our case, however, we seek for a stationary solution in the form $\psi(\xi,\tau)=g(\tau)e^{iq\xi}$ and substituting it in Eq. (3) we can have the equation for $g(\tau)$ as,
@@
$\Delta\tau_{w}=\mathcal{A}_{0}^{2}\tau_{sh}\digamma_{s}\xi,$ (8)
where a hyperbolic secant function is considered as an ansatz and the parameter ϝ$\digamma_{s}$ is defined as, ϝ$\digamma_{s}=\frac{1}{m}-\frac{2}{m}\frac{\ln(\sqrt{m}+\sqrt{m+1})}{m\sqrt{m}\sqrt{m+1}}-\left[\frac{\cos^{-1}(1+2m)}{2m}\right]^{2}$ with $m=s\mathcal{A}_{0}^{2}$, where $\mathcal{A}_{0}$ represents the input amplitude. In Fig. 2 ($e$)-($f$), the white dotted lines indicate the analytical prediction (based on Eq. (8)) of temporal shift which shows an excellent agreement with the full numerical result.
Finally, we investigate the stability of BQS in saturable media through linear stability analysis (LSA). In this approach, the stationary QS, denoted as $g(\tau)$, is subjected to small amplitude perturbations represented by $v(\tau)$ and $w(\tau)$. The optical field is expressed as $\psi(\tau,\xi)=\left[g(\tau)+w^{*}(\tau)e^{h^{*}\xi}+v(\tau)e^{h\xi}\right]e^{iq\xi}$. By substituting this form into the governing equation Eq. (3) and performing linearization with respect to the perturbations, we derive two linear systems for $v$ and $w$ that yield the growth rate $h$. This process culminates in an eigenvalue problem as follows,
@@
The stability phase plot (see Fig. 3) in the $\mathcal{A}$-$s$ parametric space reveals that BQSs are stable within a limited region where the maximum instability eigenvalue, $\Re[h_{max}]$, is zero. To confirm the LSA findings, QS stability was examined by introducing a $10\%$ noise in pulse amplitude at three distinct settings. Utilizing the Crank-Nicolson algorithm, the evolution of the pulse was analyzed for $s=0.5$ and $\delta_{4}=-0.0125$. The results indicated that the QSs were linearly unstable at points 1 ($\mathcal{A}=6$, $s=0.5$) and 3 ($\mathcal{A}=1$, $s=0.5$), whereas they exhibited stability at point 2 ($\mathcal{A}=1.78$, $s=0.5$). The linear-stability spectra revealed that points 1 and 3 contained positive real part eigenvalues, confirming linear instability, in contrast to point 2, where all eigenvalues were imaginary, indicating linear stability. Additionally, in all cases, the continuum eigenvalue edges at $\Im(h)=\pm q$ and the pair of discrete eigenvalues on the imaginary axis indicates the internal modes contributing to shape oscillations in the soliton pedestal [38].
In summary, we demonstrate the existence of novel BQS within an optical medium characterized by SN. Through a theoretical analysis employing variational optimization, we find that unique bistable solitonic states with varying amplitude profiles yet identical widths can persist. A realistic waveguide is engineered using the LN crystal, promoting an advantageous dispersion environment for BQS excitation. A perturbative variational analysis is also employed to grasp the influence of shock on BQS. Finally, a linear stability analysis is conducted to confirm the stability of this unique structure against perturbations. Our findings enhance the understanding of self-organized temporal structures as they develop under specific dispersion environment in saturable nonlinear media, indicating potential applications in high-power laser systems and communication technologies.
Acknowledgment
T.D. and S.R. was financially supported by IIT Kharagpur, the Ministry of Education of the Government of India, and A.P.L. is supported by Anusandhan National Research Foundation(ANRF) erstwhile SERB, Government of India.